### v2

# AI-enhanced Collective Intelligence: The State of the Art and Prospects

Abstract.
The current societal challenges exceed the capacity of human individual or collective effort alone. As AI evolves, its role within human collectives is poised to vary from an assistive tool to a participatory member. Humans and AI possess complementary capabilities that, when synergized, can achieve a level of collective intelligence that surpasses the collective capabilities of either humans or AI in isolation. However, the interactions in human-AI systems are inherently complex, involving intricate processes and interdependencies. This review incorporates perspectives from network science to conceptualize a multilayer representation of human-AI collective intelligence, comprising a cognition layer, a physical layer, and an information layer. Within this multilayer network, humans and AI agents exhibit varying characteristics; humans differ in diversity from surface-level to deep-level attributes, while AI agents range in degrees of functionality and anthropomorphism. The interplay among these agents shapes the overall structure and dynamics of the system. We explore how agents’ diversity and interactions influence the system’s collective intelligence. Furthermore, we present an analysis of real-world instances of AI-enhanced collective intelligence. We conclude by addressing the potential challenges in AI-enhanced collective intelligence and offer perspectives on future developments in this field.
AI, Collective Intelligence, Hybrid Intelligence, Multi-agent systems, Human-machine networks
1. Introduction
Our societies form, sustain, and function because of our intelligence. In the animal kingdom, the more intelligent the animals are, the more social they are with each other. Dunbar’s ”social brain theory” (Dunbar, 1998) even suggests that significantly more superior human intelligence compared with other primates is the result of our need to be able to manage and maintain our social lives; it is humans’ outstanding sociality that forced them to become more intelligent. The bottom line is that in human societies, collectives are intelligent, and intelligence is collective. As a concept, collective Intelligence (CI) refers to the emergent outcome of the collective efforts of many individuals. The superiority of collective intelligence to the intelligence of any of the individuals who contributed to it has been demonstrated scientifically and practically in numerous instances (Bonabeau, 2009).
In every new chapter in the history of Information and Communication Technologies (ICT), collective intelligence has been elevated to a higher level through more readily and cheaply available platforms for collaboration and exchanging ideas. The Internet, in general, and in particular, the World Wide Web, was designed to foster large-scale collaboration. Tim Berners-Lee’s initial intention in developing HTTP, which later led to the development of the WWW, was to facilitate collaboration between CERN researchers. Some thirty years later, the Web facilitates the largest collaborative projects in human history, such as Wikipedia (Yasseri and Kertész, 2013), citizen science projects (Ibrahim et al., 2021), collaborative software development, and collaborative tagging projects (Yasseri and Menczer, 2023).
Even though large-scale knowledge-generating collaborations, crowdsourcing, and, in a more general form, collective intelligence are not Internet phenomena per se, we can see this new technology has undoubtedly boosted the state-of-the-art in collaborative knowledge creation, task execution, and collective decision-making. Artificial Intelligence (AI), similar to the Web and Internet-based technologies, is reshaping different aspects of our lives at a tremendous pace. In light of this rapidly evolving situation, it is essential to consider how, similar to previous disruptions by new ICTs, AI can magnify and enhance our collective intelligence.
1.1. Collective intelligence
Intelligent agents (humans, animals, and intelligent artifacts) often interact with others to tackle complex problems. They combine their knowledge and personal information with information acquired through interaction and social information, resulting in superior collective intelligence. The concept of collective intelligence has been redefined over the years, primarily due to different research streams that have employed the term in qualitatively different ways (Woolley and Gupta, 2023).
One stream of research uses collective intelligence to refer to the outcome of collaboration produced by amalgamation of the input from a large crowd, exemplified in scenarios such as online contests or crowdsourced science (Aristeidou and Herodotou, 2020; Kittur et al., 2009; Malone, 2018b). Here, the concept of collective intelligence is often mixed up with the wisdom of the crowd. The wisdom of the crowd refers to judgments achieved by large groups and crowds consisting of distributed individuals through the pooling of personal information, often via simple averaging (Surowiecki, 2004). It relies on the statistical aggregation of independent judgments to collectively arrive at a decision or prediction that is often better than the answer given by any single member of the group. In contrast, collective intelligence is more related to achievements by smaller collaborating groups and pairs, traditionally studied within the broader topic of group performance and decision-making (Kerr and Tindale, 2004; Malone et al., 2010). A more theoretical distinction is based on the fact that collective intelligence emerges from the synergy of interactions between intelligent agents (Grasso and Convertino, 2012) collaborating towards a common goal through interconnected efforts.
Another stream of research defines collective intelligence as an ability that ”can be designed to solve a wide range of problems over time in the face of complex and changing environmental conditions” (Woolley and Gupta, 2023). Converging evidence suggests the presence of a general collective intelligence factor (”c-factor”) that serves as a predictor for a group’s performance across a wide range of tasks  (Woolley et al., 2010, 2015), similar to the individual’s general intelligence (g-factor) but extended to groups. Research has found that this ”c-factor” is not strongly correlated with the average or maximum individual intelligence of the group members, suggesting that collective intelligence is more than just the sum of the individual intelligence present in the group (Woolley et al., 2010). In this sense, this definition is virtually compatible with the one above.
A group’s collective intelligence is fundamentally influenced by its composition and interactions. Past work indicates a positive correlation between the ”c-factor” and factors such as the average social sensitivity  (Woolley et al., 2010), the equitable distribution of conversational participation  (Woolley et al., 2010), and the percentage of female members in the group  (Woolley et al., 2010). Groups with diverse perspectives and skill sets are likelier to cultivate innovative solutions that might not emerge in a more homogenous group. Effective interaction processes, encompassing coordination, cooperation (Malone and Bernstein, 2022), and communication patterns (Engel et al., 2015) are crucial. Prior studies have identified three distinct socio-cognitive transactive systems  (Woolley et al., 2023a) responsible for managing collective memory (Yasseri et al., 2022), collective attention (Gildersleve et al., 2023), and collective reasoning (Wyss and Beste, 2017), all of which are essential for the emergence of collective intelligence.
1.2. Artificial intelligence
The definition of artificial intelligence has undergone numerous evolutions in the past semi-century. Artificial intelligence can be understood as the simulation of human intelligence processes by machines (Konar, 2018), especially computer systems. Previous research considers the classification of AI into two distinct categories (Hassani et al., 2020). The first classification categorizes AI based on the level of human-like cognitive abilities they exhibit, such as the ability to ”think” and ”feel.” There are four primary AI types (Burns et al., 2021): reactive AI, limited memory AI, theory of mind AI, and self-aware AI. Among them, the limited-memory AI is the most common at the current time, ranging from virtual assistants to chatbots to self-driving vehicles. Such AI can learn from historical data to recognize patterns, generate new knowledge and understanding, and inform subsequent decisions.
The second classification is a technology-oriented approach that categorizes AI into Artificial Narrow Intelligence (ANI), Artificial General Intelligence (AGI), and Artificial Superintelligence (ASI). ANI involves creating computer systems capable of performing specific tasks in a manner similar to human intelligence (Rawat et al., 2023), but often surpass humans in efficiency and accuracy (Kuusi and Heinonen, 2022). ANI machines have a narrow range of capabilities and can represent all existing AI. AGI refers to machines that exhibit the human ability to learn, perceive, and understand a wide range of intellectual tasks (Goertzel, 2014; Review, 2023). Finally, the primary goal of ASI is to develop a machine that exhibits cognitive abilities higher than those of humans (Hassani et al., 2020).
Existing AI can perform tasks including but not limited to complex calculations, language translation, facial recognition, and financial market prediction. Recent development in generative AI, such as OpenAI’s text-to-video model Sora (OpenAI, 2024), exemplifies AI’s potential in creative industries. As AI technology evolves, it is poised to bring groundbreaking advancements across various fields. Beyond content creation, as AI deepens the understanding of the physical world and develops its simulation ability at the perceptual and cognitive level, it is possible to develop more super-intelligent tools in various other fields.
AI’s ability to process vast amounts of data on a scale is far beyond human capabilities. However, due to its limitations, many real-world applications cannot yet be solved solely by AI. AI currently lacks the deep conceptual and emotional understanding that humans possess about objects and experiences (De Cremer and Narayanan, 2023). AI cannot interpret human language’s nuances, including contextual, symbolic, and cultural meanings (Nath and Sahu, 2020; Mitchell, 2021). AI cannot be an ethical decision-maker due to a lack of human attributes of intentionality, care, and responsibility (De Cremer and Narayanan, 2023). Therefore, combining human insight with AI’s analytical power is crucial for addressing complex real-world challenges, leveraging the strengths of both to compensate for each other’s weaknesses.
1.3. Human-AI hybrid collective intelligence
Even though the field of collective intelligence initially focused solely on groups of people, in recent years, it has gradually expanded to include AI as group members in a new framework referred to as ”hybrid intelligence” (Dellermann et al., 2019). Proponents of the hybrid intelligence perspective stress that humans and AI can connect in ways that allow them to collectively act more intelligently and achieve goals that are unreachable by any of the individual entities alone (Akata et al., 2020). Researchers investigating this hybrid collective intelligence explore ”how people and computers can be connected so that collectively they act more intelligently than any person, group, or computer has ever done before” (Malone, 2018a).
Acknowledging the complementary capabilities of humans and AI as discussed in the previous section, researchers identify the need for developing socio-technological ensembles of humans and intelligent machines that possess the ability to collectively achieve superior results and continuously improve by learning from each other (Dellermann et al., 2021). Previous research provides evidence that teaming humans with AI has the potential to achieve complementary team performance (CTP), a level of performance that AI or humans cannot reach in isolation (McNeese et al., 2021b). For instance, a study demonstrates that humans can use contextual information to adjust the AI’s decision (Hemmer et al., 2022). Research on mixed teams composed of both humans and AI shows hybrid teams could achieve higher performance metrics, such as team situational awareness and team score, compared to all-human teams (McNeese et al., 2021b).
1.4. Empirical evidence and theoretical gaps
Technological advances have constantly disrupted how we produce, exchange, collect, and analyze information and consequently make decisions, individually or collectively. Automation through smart devices based on machine learning, knowledge graph-based machine reasoning, and walking and talking household devices leveraging natural language processing have already reshaped our personal, professional, and social lives. Collective intelligence cannot be an exception. Inevitably, our collective decision-making processes have been, and will be, disrupted by AI.
Machine learning and automation can increase the efficiency and scalability of collective intelligence in Citizen Science projects, where citizens volunteer to help scientists tag and classify their large-scale datasets (Jackson et al., 2018). Members of a crowd with different interests and areas of competence can be matched to other tasks by recommendation systems (Yuen et al., 2011). Generative algorithms can be used to extrapolate human solutions and generate new ideas (Scirea et al., 2015). Clustering algorithms can be used to reduce the solution space of a complex task as humans are exploring possible solutions (Hocking et al., 2018). Machines can unify similar solutions to mitigate statistical noise (Zhang et al., 2015). Matching algorithms can be used to match individuals and build efficient groups (Karger et al., 2011).
This review focuses on the integration of AI to bolster the collective intelligence of human groups, addressing a notable theoretical gap in understanding performance enhancements, particularly in hybrid human-AI configurations. With AI integration becoming pervasive across various sectors, exploring how this collaboration can unlock optimal capabilities is imperative. This exploration is essential not only for boosting productivity and fostering innovation but also for ensuring that AI complements and enhances human skills rather than replacing them outright. The industry’s pivotal role in driving current AI-CI research underscores the absence of a comprehensive theoretical framework (Berditchevskaia et al., 2022).
In subsequent sections, we introduce a dedicated framework to deepen the comprehension of AI-enhanced collective intelligence systems. We then delve into analyzing applications and discussing the implications of such fusion. Finally, we conclude by addressing existing challenges in designing AI-CI systems and speculate on the future trajectory of the field.
2. AI-enhanced collective intelligence framework
2.1. Multilayer representation of the human-AI complex system
In advancing the field of human-AI collective intelligence, prior research shows the necessity of formulating theories encompassing collective intelligence, combining both humans and AI (Eide et al., 2016). Developing such theoretical frameworks requires an in-depth comprehension and interpretation of human-AI systems, which are characterized by their high complexity and interrelated processes. Network science offers tools that enable us to understand the complexity of social systems (Pedreschi et al., 2023) where a multitude of interacting elements give rise to the collective behaviors of the whole system (Bianconi et al., 2023). To enhance our understanding of collective intelligence in the human-AI system, and explore how ”the whole is more than the sum of the individual parts” (Aristotle and Aristotle, 1933; Anderson, 1972), we integrate complexity science and network science approaches, and propose a multilayer representation of the complex system involving human and AI agents.
Figure 1. A multilayer representation to untangle the processes in the complex system of human and AI agents, with three layers influencing each other: the cognition layer, the physical layer, and the information layer. External factors and the changing environment can influence the collective intelligence of the whole system.
As shown in Fig. 1, a real-world collective intelligence system can be mapped into a multilayer network setting where three layers influence each other: cognition, physical, and information. The interacting agents are represented as nodes, and the relations are represented as links connecting the nodes in the network. The construction of such networks and the meaning of the links are context-dependent. A node can also represent a group of agents instead of an individual agent in situations where the group behaves or is treated as a single entity within the network. Across layers, the nodes are the same, but the interactions are different.
The cognition layer is only indirectly perceivable, which contains mental processes. Intelligent processes are involved during problem-solving, such as sense-making, remembrance, creation, coordination, and decision-making, often happening in the cognition layer. The links are hard to determine, but these processes exist and are fundamental to the emergence of collective intelligence in the whole system. In opposition to this indirectly perceivable layer, there is the physical layer where humans and AI have tangible physical interactions. Information interaction in the information layer refers to exchanging information between agents through various communication channels. Even though our current framework revolves around pairwise interactions between nodes, recent advancements in the study of higher-order networks, delving beyond pairwise interactions (Battiston et al., 2020), can significantly enrich our understanding of the collective emergence of intelligence.
In addition to internal processes, it is essential to acknowledge the influence of the environment on the system (Janssens et al., 2022). The human-AI complex system functions in a potentially dynamic environment. Previous research suggests that collective intelligent behavior depends on the environment (Janssens et al., 2022), as certain behaviors or strategies that work well in one setting may not be as effective or relevant in another (Kämmer et al., 2014).
Human-AI hybrid systems can be viewed as complex adaptive systems, continually evolving and adapting through interactions within dynamic environments (Goldstein, 2011). The concept of emergence in complex systems can be used to describe the phenomenon where new properties emerge at the collective level, which is not present at the level of individual components (Goldstein, 2011; Meehl and Sellars, 1956; Girvin and Yang, 2019). Here, the collective intelligence of the whole system is such an emergent property superior to the micro-level individual intelligence. This property, which includes outcomes, abilities, characteristics, and behaviors, is aligned with and encompasses the major existing definitions of collective intelligence. It emerges through complex nonlinear relationships between the agents and is likely a result of both bottom-up and top-down processes (Woolley et al., 2015). The former encompasses the aggregation of group-member characteristics that foster collaboration. The latter includes structures, norms, and routines that govern collective behavior, thereby influencing the efficacy of coordination and collaboration (Woolley et al., 2015). Employing a multilayer representation of this complex system in a changing environment can facilitate a deeper understanding of the interactions and relationships between human and AI agents, untangling the intricate processes involved in the emergence and maintenance of collective intelligence.
2.2. The goal and the task
In the previous section, we viewed the hybrid human-AI groups as complex systems that often function in dynamic and complex task environments. Addressing challenges within such systems necessitates a clear understanding of the specific tasks’ goals and nature. The system’s goal might diverge from individual objectives, highlighting the need for strategic coordination to balance collective aims with personal pursuits.
Task types can vary significantly, from generative activities such as brainstorming, which requires creative and divergent thinking, to analytical tasks such as solving Sudoku, which demands logical reasoning and pattern recognition (Delahaye, 2006). The diversity of the tasks requires different abilities, including memory utilization, creative imagination, sense-making, and critical analysis. In the following subsections, we delve into details of group-member diversity’s influence, the structure and dynamics of interaction networks, and AI’s role in enhancing group performance.
2.3. Diversity
The relationship between diversity and group collective intelligence is highly complex due to the richness of the notion of diversity and its effects under different conditions. A group can exhibit surface-level and deep-level diversities (Phillips and Loyd, 2006). The surface-level diversity comes from readily observable social categories, including gender, age, and ethnicity (Wegge and Schmidt, 2009). The deep-level diversity refers to differences in psychological characteristics (Harrison et al., 2002), and can include personality, cognitive thinking styles, and values.
Diversity has shown benefits for work teams in terms of more creativity, improved decision-making, and access to a broader pool of talent  (Bagshaw, 2004; Van der Vegt and Janssen, 2003). However, diversity can also have adverse effects, such as emotional conflict (Pelled et al., 1999; Yasseri et al., 2012b), stress (DeFrank and Ivancevich, 1998), poor work relationships, and poor overall performance (Van Knippenberg et al., 2004). There is a lack of consensus on which aspects of group diversity are likely to result in positive outcomes and which aspects could produce negative outcomes (Van Knippenberg et al., 2004; Jansen and Searle, 2021). The effects of diversity on group performance depend on the complexity of the task and various moderators, such as the density of the interaction network (Baumann et al., 2024).
2.3.1. Surface-level diversity
Gender
The evidence concerning the effect of gender diversity on team performance is equivocal and contingent upon various contextual factors (Bear and Woolley, 2011). Past research found that gender-diverse teams outperform gender-homogeneous teams when perceived time pressure is low (Kearney et al., 2022). One study showed gender diversity was negatively associated with performance, but only in large groups (Wegge and Schmidt, 2009). Previous evidence strongly suggests that team collaboration is greatly improved by the presence of females in the group (Bear and Woolley, 2011; Woolley et al., 2010). Gender-diverse teams produce more novel and higher-impact scientific ideas (Yang et al., 2022). Other studies show that gender composition seems to matter for team performance. However, when controlling for the individual abilities of team members, the relation between gender composition and team performance vanishes (Städter et al., 2022). One study points out that the ”romance” of working together can benefit group performance (Zhang and Hou, 2012). Gender-diverse groups perform better than homogeneous groups by decreasing relationship and task conflict (Zhang and Hou, 2012).
Age
Age diversity may potentially enhance group performance, but the benefits may depend on the context. Several studies have found that age diversity in teams positively correlated with team performance when completing complex tasks (Wegge et al., 2012).
Past research points out the positive effects were only seen under conditions of positive team climate and low age discrimination (Wegge et al., 2012). Another study shows teams with more age diversity reported more age discrimination (Kunze et al., 2013), associated with lower commitment and worse performance.
Ethnicity
There are inconsistent results in the case of ethnic diversity. One study (McLeod et al., 1996) found racially diverse groups produced significantly more feasible and effective ideas than homogeneous groups. Another study found racial similarity in groups associated with higher self-rated productivity and commitment (Riordan and Shore, 1997). Under the right conditions, teams may benefit from diversity in ethnicity and nationality (Nielsen et al., 2017).
AI exhibits surface-level diversity through different anthropomorphic features in robots, such as voice, avatars, and human-like characteristics. Research has indicated that the gendering of AI can affect perceptions and trust levels in users. For instance, studies have found that users may exhibit more trust in AI personal assistants whose voice gender matches their own (Lee et al., 2021). Assigning a female gender to AI bots enhances their perceived humanness and acceptance (Borau et al., 2021). Previous studies have indicated a perception bias in AI agents based on gender: male AI agents tend to be viewed as more competent, whereas female AI agents are often perceived as warmer (Ahn et al., 2022). The perceived gender of the machine can make the social dynamics in hybrid teams even more complicated. However, these gendered characteristics in AI also raise concerns about reinforcing stereotypes and biases (Chin and Robison, 2020). The influence of AI’s gendered traits on human interactions is an ongoing area of research, with implications for how AI is designed and utilized.
2.3.2. Deep-level diversity
Personality
Some work proposes that investigations into the relationship between personality and work-related behaviors should expand beyond the linearity assumptions (Curşeu et al., 2019), showing that extraversion, agreeableness, and conscientiousness have inverted U-shaped relationships with peer-rated contributions to teamwork (Curşeu et al., 2019). Similar work demonstrates a curvilinear relationship between a team’s average proactive personality and performance (Zhang et al., 2021a). This body of work also identifies the moderating effects of personality dispersion, team potency, and cohesion on this relationship (Zhang et al., 2021a). Previous research has found that a team’s collective openness to experience and emotional stability moderates how task conflict affects team performance (Bradley et al., 2013). Specifically, in teams with high openness or emotional stability, task conflict positively influences performance (Bradley et al., 2013).
Cognitive style
Team cognitive style diversity refers to the variation in team members’ ways of encoding, organizing, and processing information (Aggarwal et al., 2019). Research indicates its significance in fostering innovation within teams (Aggarwal and Woolley, 2019). A certain level of cognitive diversity contributes to collective intelligence by providing diverse cognitive inputs and viewpoints necessary for task work (Aggarwal and Woolley, 2019). However, excessive diversity may lead to high coordination costs (Straub et al., 2023), as team members with differing perspectives struggle to understand each other (Cronin and Weingart, 2007; Mello and Rentsch, 2015). Studies suggest that the relationship between cognitive style diversity and collective intelligence is curvilinear, forming an inverted U-shape (Aggarwal et al., 2019). Additionally, the beneficial effects of cognitive diversity might be mediated through factors such as task reflexivity and relationship conflict (Chen et al., 2019).
Value judgment
Values are internalized beliefs that can guide behavior and enhance motivation (Cennamo and Gardner, 2008). Past research found higher diversity in values was associated with more conflict (Liang et al., 2012) in relationships, tasks, and processes. Overall, existing research consistently indicates that value congruence within teams positively influences performance, satisfaction, and conflict (Jansen and Searle, 2021) and moderates the effects of informational diversity (Jehn et al., 1999).
While AI systems can display various behaviors or responses based on their programming and training, this diversity is a product of different algorithms, models, and data inputs. Unlike humans, AI does not have personal experiences, beliefs, or innate personality traits that contribute to deep-level diversity. In contrast, AI might influence deep-level diversity in human teams, primarily by shaping social interactions and decision-making processes. The optimal composition of teams remains uncertain (Bell, 2007). A previous review underscores the need for in-depth research into both surface and deep-level diversity within teams (Jansen and Searle, 2021). This research should explore how these aspects of diversity influence various outcomes, examining their effects over time and across different levels in a broad array of contexts (Jansen and Searle, 2021).
2.4. Interactions
Interaction distinguishes a group from a mere collection of individuals (Janssens et al., 2022): one person’s behavior forms the basis for the responses of others (Driskell and Salas, 1992). Therefore, the interaction between two members establishes a link in the interaction network. In terms of the group’s interaction network structure, past research suggests that the small-worldness of a collaboration network (small network diameter and high clustering) improves performance up to a certain point, after which the effect will reverse (Uzzi and Spiro, 2005). An ”inverted U” relationship is found between connectedness and performance (Lazer and Friedman, 2007), suggesting there is an optimum number of connections for any given size of the collaboration network. A modular structure (a large number of connections within small groups that are loosely connected) can increase efficiency (Navajas et al., 2018). A study on forbidden triads among jazz groups shows that heavily-sided open triangles are associated with lower success (measured by the number of group releases) (Vedres, 2017).
AI can contribute to human groups in various ways by augmenting existing human skills or complementing capabilities that humans do not possess. AI has attributes beyond humans, such as more extended memory, higher computational speed, and a more vital ability to work with large, diverse datasets. AI’s augmentation of human cognition enables teams to navigate complex situations effectively (Woolley et al., 2023b). Additionally, AI may also help reduce the harm of implicit bias in humans (Lin et al., 2021) and shape better decisions. Taking advantage of the differences between humans and AI can contribute to the performance of hybrid human-AI groups. In the following sections, we discuss the rules and incentives that govern the interaction processes and the structure and dynamics of the group’s interaction network.
2.4.1. Rules and incentives
Decision rules function as prescribed norms that direct interactions and play a crucial role in shaping the communication and information integration within a group (Meslec et al., 2014). Previous research has focused on disentangling decision rules guiding the team’s interaction, ultimately fostering synergy. Some literature emphasizes that the decision rules themselves are intelligent (Wolf et al., 2015). In addition to the rules guiding behavior, the group members interact with each other under specific incentive schemes. The intrinsic and extrinsic incentives are the driving forces (Giacomelli, [n. d.]) for the agents to behave in a direction such that the social network becomes dynamic and moves towards a collective goal. Some research proposes that an incentive scheme rewarding accurate predictions by a minority can foster optimal diversity and collective predictive accuracy (Mann and Helbing, 2017).
AI can augment human cognition (Regens, 2019) to help teams adapt to complexity. While AI can simulate certain human-like incentives through programming and learning algorithms (Park et al., 2023), fully capturing the depth and nuance of human behavior, such as irrationality or altruism, is challenging. A recent study applying a Turing test to compare AI chatbots and humans found that AI chatbots exhibited more altruistic and cooperative behavior than their human counterparts in behavioral games (Mei et al., 2024). Machines can be programmed to mimic human behaviors, but whether they genuinely ”exhibit” them like humans is debatable. In the context of collective intelligence, integrating such human-like behaviors in AI may enhance collaboration and decision-making within diverse teams. However, the extent to which AI can be truly self-aware and authentically replicate complex human traits remains a subject of ongoing research and philosophical debate.
2.4.2. Group structure and dynamics
Group size
Much research on collective intelligence has been dedicated to finding a group’s optimum size and structure. Several studies have reported that performance improves with group size through enhanced diversity (Malone and Bernstein, 2022), whereas others suggest the opposite: small groups are more efficient (Galesic et al., 2018). Another study suggests that large teams develop and small teams innovate (Wu et al., 2019). Larger groups can better utilize diverse perspectives and knowledge to solve complex problems but also tend to experience more coordination problems and communication difficulties (Woolley et al., 2010). The consideration of optimal group size can be related to the dynamics of the group, whether it is static, growing, or diminishing.
In human-AI hybrid teams, the concept of group size becomes nuanced with AI integration, especially when AI entities such as Large Language Models (LLMs) are involved. The countability of AI agents, particularly if multiple agents are derived from a single LLM, raises questions about their distinctiveness and individuality in a team context. Unlike humans, AI agents can simultaneously process multiple tasks, blurring the traditional boundaries of group size. Applying past research on human group dynamics to these hybrid systems requires rethinking the notions of individual contribution, team cohesion, and communication. It is important to consider how AI’s unique capabilities and scale impact these dynamics and whether multiple AI agents from a single model represent distinct entities or a collective intelligence resource.
Structure
research shows that network structure, specifically hierarchy, and link intensity, can affect information distribution (Tang and Liao, 2023), influencing group performance. Another study observed that the presence of structural holes (Burt, 2018) in leaders’ networks and the adoption of core-periphery and hierarchical structures in groups correlated negatively with their performance (Cummings and Cross, 2003). Incorporating the temporal dimension, adaptive social networks can significantly impact collective intelligence by allowing the network’s structure to evolve based on feedback from its members (Almaatouq et al., 2020).
Team size moderates the relationship between network structure and performance (Yuan and Van Knippenberg, 2022). Studies have found central members can coordinate with other team members more easily in smaller teams (Balkundi and Harrison, 2006), whereas in larger teams, communication challenges may arise (Hollenbeck et al., 2002). A recent study shows a more central connected leader in advice-giving networks has a more positive impact on the performance of larger teams, but this effect is reversed in smaller teams (Yuan and Van Knippenberg, 2022).
Determining the optimal size, structure, and human-AI ratio for hybrid teams is a complex task requiring further study, particularly concerning task-specific requirements. Moreover, incorporating AI into human teams can reshape group hierarchies, norms, and rules, thereby affecting collective intelligence.
Communication patterns
Teams often outperform individuals primarily due to explicit communication and feedback within the team (Larson Jr, 2013). Understanding the communication network in crowds is crucial for effectively designing and managing crowdsourcing tasks (Yin et al., 2016). Various studies highlight that centralized and decentralized communication patterns can effectively promote team performance, contingent on the nature of the task and the team’s composition (Cohen et al., 1962; Rosen, 2008). In particular, teams handling complex tasks tend to be more productive with decentralized communication networks (Shaw, 1964). Conversely, the centralization of communication around socially dominant or less reflective individuals can negatively impact the utilization of expertise and overall team performance (Sherf et al., 2018). Communication patterns, such as equal participation and turn-taking of team members, can positively affect collective intelligence (Woolley et al., 2010).
The integration of AI in human groups can influence the conversation dynamics in different ways, depending on the role of AI. AI chat interventions can improve online political conversations at scale (Argyle et al., 2023), reducing the chance of conflicts. While AI machines might struggle with capturing the subtle and ineffable social expressions that make up the dynamics of human groups, this can be advantageous. In certain contexts, AI can be social catalysts (Rahwan et al., 2020) to promote communication where human capacity is limited. Striking the right balance in AI’s role is key to maintaining the natural dynamics of human interaction.
Cognition processes
Cognitive processes are foundations for the emergence of collective intelligence (Woolley and Gupta, 2023), encompassing knowledge acquisition, information processing, problem-solving, decision-making, language, perception, memory, attention, and reasoning. Human cognition processes evolve and interact with each other during various tasks and experiences. Humans can represent the mental states of others, including desires, beliefs, and intentions (Rabinowitz et al., 2018), known as the Theory of Mind (ToM). Past research, inspired by ToM, showed that the ability of team members’ to assess the mental state of others is positively associated with team performance both in face-to-face settings (Woolley et al., 2010) and online (Engel et al., 2014). Regarding collective understanding within a group, the Shared Mental Models (SMM) theory is helpful for understanding, predicting, and improving performance in human teams (Andrews et al., 2023).
Applying theories such as SMM to human-AI teams is a natural progression in research. These models can enhance team performance by ensuring that human and AI members have a similar understanding of tasks and each other (Andrews et al., 2023), facilitating better prediction of needs and behaviors. Developing ToM in AI also improves coordination and adaptability in human-AI teams (Baker et al., 2011). This involves AI’s capability to anticipate and respond to new information and behaviors. Moreover, recent studies indicate that incorporating ”hot” cognition, influenced by emotional states in AI, can further improve human-machine interactions (Cuzzolin et al., 2020). Understanding the transactive systems (Woolley et al., 2023a) of memory, attention, and reasoning within these teams is key to grasping the members’ knowledge and task preferences.
2.5. AI modes of contribution
AI’s role in hybrid groups varies based on its autonomous agency and functionality. AI may serve as a technical tool for human assistance or as an active agent that interacts with and influences humans. Functionally, AI’s role can range from an assistant to a teammate, coach, or manager. Furthermore, the degree of AI’s anthropomorphism, from non-physical voice-only interfaces to avatars and physically embodied robots, also significantly affects its role in the group.
2.5.1. AI as an assistant
Assistant-type AI systems are generally used as technical tools that complement or augment human abilities to increase human efficiency in performing tasks. For example, language translation AI can assist in translating text or spoken language, facilitating communication across language barriers. Generative AI can inspire human creators and provide them with new ideas (Doshi and Hauser, 2023). Administrative AI assistants can help schedule meetings, manage emails, and organize tasks. Smart home AI assistants such as Alexa and Siri can help manage smart home devices, play music, provide weather updates, answer questions, and assist with daily routines. AI assistants are technical tools that help humans to coordinate tasks in an efficient manner, optimize their decision making, as well as personalize their experiences.
2.5.2. AI as a teammate
AI can function as teammates, working alongside humans with complementary abilities. AI teammates are already used in real-world settings; most employees using AI already see it as a coworker (Boston Consulting Group, 2022). In healthcare, radiologists and AI work together to diagnose from the pictures of the patients. In creative industries, AI collaborators in music, art, and literature, such as Google’s Magenta (TensorFlow, 2024) project, collaborate with musicians and artists to create new compositions or artworks using AI algorithms. Furthermore, AI robots can also serve as teammates in various workplaces, which we will discuss further in the AI embodiment section.
2.5.3. AI as a coach
As a coach, AI can provide guidance, support, and personalized feedback to individuals seeking to improve their skills and achieve their goals. In team settings, one perspective highlights the AI coach’s ability to provide a comprehensive, global view of the team’s environment, guiding players who have only partial views (Liu et al., 2021). This approach involves the AI coach strategizing and coordinating, distributing tailored strategies to each team member based on their unique perspectives and roles. Another dimension of AI coaching focuses on assessing and improving teamwork by closely monitoring team members during collaborative tasks (Seo et al., 2021). Here, the AI coach actively intervenes with timely suggestions based on inferred mental model misalignment (Seo et al., 2021) of the team members.
2.5.4. AI as a manager
AI has been used in working environments with supervision or managerial roles, such as making decisions of hiring, promotion, and reassigning tasks (Kiron et al., 2023). An AI manager may mitigate human biases (Hofeditz et al., 2022) in hiring and promotions, contributing to a more diverse workforce, which is a key aspect of collective intelligence. Additionally, by analyzing individual strengths and weaknesses and efficiently allocating tasks, an AI manager can optimize team dynamics and workflows, improving overall group performance.
2.5.5. AI embodiment
With the advancement of AI in areas such as natural language processing, object recognition, and creative idea generation, there’s a trend to integrate AI with robotics, creating physical entities in the real world. Embodied AI (Pfeifer and Iida, 2004) refers to intelligent agents interacting with their environment through a physical body. These AI-enabled machines, embodied in robotic arms or autonomous vehicles, can augment the capabilities of human workers. They are equipped with advanced sensors, motors, and actuators, enabling recognition of people and objects and safe collaboration with humans in diverse settings, including factories, warehouses, and laboratories. Furthermore, advanced AI-robotics integration allows robots to recognize and respond to human speech and gestures, impacting conversational dynamics within mixed human-robot environments. Research has shown that a vulnerable robot’s social behavior positively shapes the conversation dynamics among human participants in a human-robot group (Traeger et al., 2020). In scenarios where conflicts emerge within a human group or in groups of strangers, a humorous robot can potentially serve as an ice breaker to intervene and resolve the awkward situation (Press and Erel, 2023), demonstrating the potential of robots to influence the nature of human-human interactions notably.
2.6. Human-AI collective intelligence emerging factors
2.6.1. Perception and reaction
In addition to investigating how humans perceive individual AI, previous research suggests the necessity of additional work to explore human perceptions of AI in collaborative team environments (Kocielnik et al., 2019). For human-AI teams to succeed, human team members must be receptive to their new AI counterparts (Harris-Watson et al., 2023). People’s pre-existing attitudes toward AI were found to be significantly related to their willingness to be involved in human-AI teams (Zhang et al., 2021b). Research finds that two aspects of social perception, warmth, and competence, are critical predictors of human receptivity to a new AI teammate. Psychological acceptance is positively related to perceived human-AI team viability (Harris-Watson et al., 2023). Regarding human perception of AI behaviors, one study finds that people cooperate less with benevolent AI agents than with benevolent humans (Karpus et al., 2021).
Humans tend to be more accepting of AI when it makes mistakes. Some studies find that users have a more favorable impression of the imperfect robot than the perfect robot when the robot behaves adequately after making mistakes (Yasuda and Matsumoto, 2013). When both human drivers and AI make errors in cases of human–machine shared-control vehicles, the blame attributed to the machine is reduced (Awad et al., 2020). Research has found that errors occasionally performed by a humanoid robot can increase its perceived human likeness and likability (Salem et al., 2013). A recent study shows humans accept their AI teammate’s decision less often when they are deceived about the identity of the AI as another human (Zhang et al., 2023a). The nature of the task requested by the robot, e.g., whether its effects are revocable as opposed to irrevocable, significantly impacts participants’ willingness to follow its instructions (Salem et al., 2015). When taking AI’s advice, humans must know the AI system’s error boundary and decide when to accept or override AI’s recommendation (Bansal et al., 2019).
2.6.2. Trust
In AI-assisted decision-making, where the individual strengths of humans and AI optimize the joint decision outcome, the key to success is appropriately calibrating human trust (Zhang et al., 2020). Trust between humans and AI plays a pivotal role in realizing collective intelligence (Gupta et al., 2023). Previous research highlights the necessity of understanding factors that enable or hinder the formation, maintenance, and repair of trust in human-AI collaborations (Gupta et al., 2023). Key factors influencing this trust include the perceived competence, benevolence, and integrity (McAllister, 1995) of the AI systems, mirroring the trust dynamics in human relationships.
The level of anthropomorphism of machines can be a predictor of the trust level of the human participants (Natarajan and Gombolay, 2020). Studies indicate that while anthropomorphic features in AI agents initially create positive impressions, this effect is often short-lived (Gupta et al., 2023). Trust tends to decline, especially when the agent’s capabilities are not clearly presented, and its performance fails to meet users’ expectations (Glikson and Woolley, 2020).
Trust in a teammate may evolve over time irrespective of team performance (McNeese et al., 2021a). Knowing when to trust the AI allows the human expert to appropriately apply their knowledge, improving decision outcomes in cases where the model is likely to perform poorly (Zhang et al., 2020). A previous study also proposes a framework highlighting trust relationships within human-AI teams. It acknowledges the multilevel nature of team trust, which considers individual, dyads, and team levels as a whole (Ulfert et al., 2023). Another study points out that the trust level of a team needs to be uniform, as uncooperative members could undermine the team’s ability to reach intended goals (De Jong et al., 2021).
2.6.3. Explainability
In human-AI decision-making scenarios, people’s awareness of how AI works and its outcomes is critical in building a relationship with the system. AI explanation can support people in justifying their decisions (Ferreira and Monteiro, 2021). A study on content moderation found that when explanations were provided for why content was taken down, removal decisions made jointly by humans and AI were perceived as more trustworthy and acceptable than the same decisions made by humans alone (Wang and Kim, 2023). Many researchers have proposed using explainable AI (XAI) to enable humans to rely on AI advice appropriately and reach complementary team performance (Hemmer et al., 2022).
Previous research finds that AI explanations can increase human acceptance of AI recommendations, regardless of their correctness (Bansal et al., 2021), leading to an over-reliance, which threatens the performance of human-AI decision-making. Whereas another study finds that there are scenarios where AI explanations can reduce over-reliance (Vasconcelos et al., 2023) using a cost-benefit framework. Some studies show that cognitive forcing can reduce over-reliance compared to the simple explainable AI approaches (Buçinca et al., 2021), with a trade-off that the designs most effective in reducing over-reliance were the least favored by users (Buçinca et al., 2021). The impact of AI explanations on decision-making tasks is contingent upon individuals’ diverse levels of domain expertise (Wang and Yin, 2021).
3. Applications and implications
The AI-enhanced collective intelligence approach is increasingly applied across various domains, offering innovative solutions to challenges such as community response to climate change, environment, and sustainability challenges (Nesta, 2024), etc. AI serves as a vital decision support tool for policymakers in detecting misinformation (Kou et al., 2022) and real-time crisis management, exemplified during the COVID-19 pandemic. Its application extends to high-stake areas, including medical diagnosis (Lebovitz et al., 2022) and criminal justice (Rigano, 2019). AI’s integration into collective intelligence efforts enables significant amplification of its impact and scalability (Verhulst, 2018).
Figure 2. Distribution of AI-enhanced collective intelligence cases by application area based on dataset curated by Supermind Design (Design, 2024).
Here, we present our analysis utilizing the Supermind Design database (Design, [n. d.]), which includes 857 categorized cases of AI-enhanced collective intelligence by application areas. Notably, while some cases might not directly involve AI or collective intelligence depending on the exact definitions, the database collectively offers a comprehensive overview. It provides insights into the state of the art in AI-CI applications. As shown in Fig. 2, the majority of applications, approximately twenty percent, are found in the public sector and NGOs, followed by high-tech, media, telecommunications, entertainment, and hospitality. Fewer instances are found in supply chain, real estate, and agriculture. It is important to note that several cases classified under ”Public sector, NGO” are private sector initiatives addressing public issues. We will analyze and highlight specific AI-CI examples across these diverse domains in the upcoming sections.
3.1. Public sector, NGO
Safecity (Safecity, [n. d.]) is a platform that crowdsources personal stories of sexual harassment and abuse in public spaces. It collaborates with the AI partner OMDENA (Omdena, [n. d.]) to build solutions to help women at risk of sexual abuse by predicting places at a high risk of sexual harassment incidents utilizing convolutional neural networks and LSTM (Long-Short-Term-Memory).
Ushahidi (Ushahidi, [n. d.]) is a non-profit online platform for crowdsourcing information, mapping, and visualizing data to respond to and inform decisions on social issues and crises. It aggregates crowdsourced information and reports from individuals across the globe. Ushahidi is developing new tools to enhance data management and analysis, such as leveraging machine learning techniques to improve the efficiency of processing crisis reports, thus enabling faster response time (Nesta, [n. d.]).
While NGOs and the public sector might seem like fertile grounds for AI-CI application, it is crucial to consider the potential high risk issues that may arise and their impact during the design and deployment of such AI-CI systems. A notable instance of a data-driven tool for social good that diverged from its original accuracy and effect is the infamous case of Google Flu (Lazer et al., 2014).
3.2. High Tech (software)
Bluenove (Bluenove, 2023), a technology and consulting company that pioneers massive collective intelligence, focuses on mobilizing communities on a large scale. Their Assembl (Bluenove, 2024) platform facilitates open conversations among participants, employing advanced AI technologies such as natural language recognition, semantic analysis, and emotion analysis to categorize content published on the platform automatically.
Figure Eight (Eight, [n. d.]) (formerly known as CrowdFlower (CrowdFlower, 2024)) was a human-in-the-loop machine learning and AI company acquired by Appen (Appen, 2024) in 2019. It leverages crowdsourced human intelligence to perform tasks such as text transcription or image annotation to train machine learning algorithms, which can be used in various applications.
The recent advancements in high-tech technologies, especially the burgeoning field of LLM software, hold the potential for revolutionary developments. Nevertheless, the high risk associated with errors in automated or AI-generated software (Zhang et al., 2023b) necessitates the inclusion of human intelligence in these projects.
3.3. Media, telecommunication, entertainment, hospitality
Civil War Photo Sleuth (Sleuth, [n. d.]) combines facial recognition technology and community effort to uncover lost identities in photographs from the American Civil War era. It leverages a dedicated community with a keen interest in the US Civil War and focuses on more accurately tagging and identifying individuals in historical photographs.
Bellingcat (Bellingcat, 2024) is an independent investigative online journalism community specializing in fact-checking and open-source intelligence. It brings together a collective of researchers, investigators, and citizen journalists to publish investigation findings. Utilizing various AI tools in investigative work, Bellingcat obtains insights from massive open-source data, such as satellite imagery, photographs, video recordings, and social media posts.
Facing the emerging challenge of deepfake videos, research indicates that a hybrid system combining human judgment with AI models outperforms either humans or models alone in detecting deepfakes (Groh et al., 2022), underscoring the value of integrating human perceptual skills with AI technology in addressing the complexities posed by deepfake content.
3.4. Energy, natural resources
Litterati (de Leon, 2019) is a mobile app and community platform that combats litter by empowering people to take action. The app allows the worldwide community to contribute their observations and findings through photographing and documenting occurrences, creating a crowdsourced database of litter data. It also provides a platform for individuals to connect, share, and support each other’s initiatives. The platform uses AI, such as machine learning and computer vision techniques, to automatically recognize different objects and materials and efficiently organize the collected litter data. Similarly, OpenLitterMap (OpenLitterMap, 2023) is another example of making a clearer planet using AI-enhanced collective intelligence.
eBird (Sullivan et al., 2014) is a mobile app and community platform for biodiversity and supports conservation initiatives to protect bird species and their habitats. It harnesses the power of a global community of birdwatchers and citizen scientists to collectively report bird observations and contribute to a shared knowledge base. eBird uses computer vision techniques to identify bird species from submitted photographs and recordings and employs machine learning predictive models to forecast bird migration patterns and species distributions.
Tackling environmental challenges is crucial, and large-scale, crowd-based AI-enhanced collective intelligence projects play a significant role in this endeavor. Nevertheless, integrating AI with human efforts might dehumanize the work and demotivate human participants (Trouille et al., 2019). This balance between efficiently leveraging AI and maintaining human engagement is a key consideration in such projects.
3.5. Education and academia
Kialo Edu (Kialo, 2024) is an argument mapping and debate site dedicated to facilitating collaborative and critical thinking. Kialo uses natural language processing techniques to analyze arguments, cluster topics, and identify discussion patterns. It allows users to participate in structured debates and discussions, promoting constructive engagement among students. This interaction helps them deeply understand and analyze the core aspects of the topics under discussion.
Zooniverse (Zooniverse, 2024) is a platform for people-powered research where volunteers assist professional researchers. Galaxy Zoo is a prominent project within the Zooniverse platform. This crowdsourced astronomy project invites people to assist in the morphological classification of large numbers of galaxies. It uses AI algorithms to learn patterns and characteristics indicative of different galaxy types by training on large datasets previously classified by citizen scientists. The combination of efforts from both the crowd and AI enables Galaxy Zoo to harness the power of human expertise and the efficiency of AI to scale up the analysis and understanding of galaxies.
A study analyzing the contributor demographics of the Zooniverse project revealed uneven geographical distribution and a gender imbalance, with approximately 30% of the citizen scientists being female (Ibrahim et al., 2021). This finding highlights concerns about representativeness and diversity in citizen science initiatives, indicating potential biases in the data and insights from such projects. AI can be used to train human contributors (Zevin et al., 2017), to de-bias collective intelligence, and to evaluate and combine human solutions (Kleinberg et al., 2018).
3.6. Healthcare
The Human Diagnosis Project (Human Dx) (Dx, 2024) is an open online system of collective medical insights that provides healthcare support to patients. Human Dx enables healthcare professionals worldwide to contribute expertise and collaborate on complex diagnostic cases (Bresnick, 2017). The platform facilitates idea exchange and mutual learning to ultimately benefit patient care by utilizing machine learning techniques that automatically learn from classifying patterns in crowdsourced data. Human Dx makes the most of limited medical resources by harnessing the collective intelligence of human physicians and AI, enabling more accurate, affordable, and accessible care for those in need.
CrowdEEG (CrowdEEG, 2024) is a collaborative annotation tool for medical time series data. The project combines human and machine intelligence for scalable and accurate human clinical EEG data analysis. It trains machine learning algorithms using feedback from clinical experts and non-expert crowds to perform feature detection or classification tasks on medical time series data.
In healthcare, the use of AI-CI poses the risk of amplifying existing biases in human-only diagnosis once we combine AI with human judgment (Groh et al., 2024). Consequently, errors resulting from such biased decision-making can erode patients’ trust not only in the AI systems but also in the medical professionals utilizing them.
3.7. Financial services
Numerai (NUMERAI, 2024) is an AI-run quant hedge fund built on crowdsourced machine learning models. Numerai hosts an innovative data science tournament in which a global network of data scientists develops machine-learning models to predict stock markets. Based on the combined knowledge of the participants, Numerai combines and aggregates these crowdsourced predictive models into an ensemble model to derive investment strategies in the financial markets.
CryptoSwarm AI (AI, 2024) is a forecasting service that provides rigorous insights and intelligence on cryptocurrencies and other Web3 assets. A combination of AI technology and real-time human insights powers it. CryptoSwarm AI uses Swarm AI (AI, [n. d.]) to amplify the collective intelligence of the online communities quickly, enabling networked groups to converge on optimized solutions in real time.
In financial markets, research generally indicates positive impacts on market efficiency due to the superior information processing and optimization capabilities of machines, coupled with human anticipation of these effects. However, this improvement is sometimes disrupted by rapid price spikes and crashes, resulting from the herding behavior of machines and the challenges humans face in intervening at extremely fast timescales (Tsvetkova et al., 2024).
3.8. Supply chain, real estate
MarineTraffic (MarineTraffic, 2024a) is a web-based platform that creates a global network of vessel tracking information, enabling users to track vessels, monitor maritime traffic, analyze trends, and make informed decisions. It relies on the collective contributions of the maritime community, who voluntarily transmit real-time information about vessel positions and movements from various sources, including AIS (automatic identification system) data and satellite data. MarineTraffic employs neural network architectures for automatic maritime object detection using satellite imagery (Bereta et al., 2020). More examples of AI applications are available at MarineTraffic Research (MarineTraffic, 2024b).
Waze (Waze, [n. d.]) is a GPS navigation software that works on smartphones and other computers. It analyzes crowdsourced GPS data and user-submitted information along the route in real-time and uses community editing to ensure the accuracy of the map data. Waze uses AI algorithms to predict traffic, optimize routes (Link, 2019), and provide personalized recommendations by analyzing driver behaviors. Waze has recently implemented integration with Google Assistant, enabling users to use the application by utilizing the voice command ”Hey Google” (Google, [n. d.]).
By improving efficiency, reducing waste, and enabling more equitable distribution and accessibility of resources and spaces, AI can play a significant role in addressing inequalities in access to resources, paving the way for sustainable and efficient infrastructure development (Cowls et al., 2021).
3.9. Agriculture
WeFarm (Nesta, 2019) is a social networking platform connecting the small-scale farming community. It helps millions of African farmers meet, exchange solutions to their questions, and trade equipment and supplies. Farmers ask each other questions about agriculture and promptly receive content and ideas from other farmers worldwide through crowdsourcing. Natural language processing processes and understands the text-based messages farmers send in different regions. Machine-learning matching algorithms consider farmers’ needs and expertise, identify the most suitable responses, and present them to the farmer who posed the question.
Mercy Corps AgriFin (AgriFin, 2023) harnesses the power of AI and a global network of partners to transform agriculture for smallholder farmers. By utilizing state-of-the-art imagery, modeling, and analysis, AI enables precise management of crops and adaptation strategies against climate change impacts. AgriFin’s integration of digital technology and data, supported by a global network, empowers these farmers in a digitally interconnected world.
In light of the global food crisis and the quest for sustainable solutions, the importance of AI-enhanced collective intelligence platforms becomes increasingly critical. These platforms can significantly contribute to addressing agricultural challenges and transitioning towards global sustainable food systems (Marvin et al., 2022).
4. Challenges and outlooks
The application areas of AI-enhanced collective intelligence continually expand as technology advances and new opportunities emerge. As introduced in the previous section, successful examples of AI-enhanced collective intelligence exist in various domains. However, challenges coexist with opportunities. When realizing human-AI collective intelligence, many aspects must be considered, such as human-machine communication, trust, crowd retention, technology design, and ethical issues (Tsvetkova et al., 2017b). Here, we briefly discuss some of these challenges and call for further work on each aspect.
4.1. Communication
Clear and effective communication between humans and AI is the engine for collective intelligence to emerge (Gupta et al., 2023). Communication refers to the process of exchanging information between teammates (Salas et al., 2009). It is essential for team performance as it contributes to the development and maintenance of shared mental models and the successful execution of necessary team processes (Salas et al., 2009; Stowers et al., 2021).
For effective communication in human-AI teamwork, AI must first be able to model human information comprehension accurately and then effectively communicate in a manner understandable to humans. Current AI technologies face limitations in verbal and contextual understanding (Stowers et al., 2021) and cognitive capabilities, such as reasoning about others’ mental states and intentions (Matthews et al., 2021). Detection, interpretation, and reasoning about social cues from a human perspective is imperative to ensure effective coordination but has yet to be achieved by AI. Although, recent developments in LLMs appear promising, further advancing our understanding of how LLMs can be utilized in such a way is crucial (Lu et al., 2023; Schaeffer et al., 2024).
AI’s inability to interpret nonverbal cues and limited self-explanation hinder human collaboration. Addressing these challenges, research suggests the development of AI agents with cognitive architectures that can facilitate both a Machine Theory of the Human Mind (MToHM) and a Human Theory of the Machine Mind (HToMM) will be especially important for supporting the emergence of collective intelligence (Gupta et al., 2023).
4.2. Trust
In team success, a human member’s trust in a machine is pivotal (Nam and Lyons, 2020). Research indicates that users interact differently with AI than humans, often exhibiting more openness and self-disclosure with human partners (Mou and Xu, 2017). Developing explanatory methods that foster appropriate trust in AI is challenging yet crucial for enhancing performance in human-AI teams (Bansal et al., 2021). Meanwhile, over-reliance on AI, especially when faced with incorrect AI advice, can be detrimental and lead to human skill degradation. Hence, it is essential for humans to judiciously rely on AI when appropriate and exercise self-reliance in the face of inaccurate AI guidance (Schemmer et al., 2022), especially in high-stakes situations. More tech literacy, public understanding of AI, and advancements in the technology itself will help in overcoming these challenges.
4.3. Crowd retention
Albeit beneficial in terms of efficiency, employing AI and citizen scientists together has been reported to damage the retention of citizen scientists (Bowyer et al., 2015; Trouille et al., 2019). There is a trade-off between efficiency and retention of the crowds when deploying AI in citizen science projects, and ultimately, if all the crowds leave the projects, the performance of AI alone would decline again. One of the factors contributing to the retention of volunteers is reported to be the social ties they make in their teams (Hristova et al., 2013). The challenge of forming social bonds with AI could increase human feelings of loneliness and isolation, potentially affecting mental health. Questions about identity and recognition could arise when humans are rewarded in conjunction with AI. Studies highlight concerns regarding the anthropomorphic appearance of social robots potentially threatening human distinctiveness (Ferrari et al., 2016). Additionally, collaboration with AI may demotivate humans due to a lack of competitive drive and an over-reliance on AI, which could diminish human participation and initiative. Understanding how to balance AI collaboration with maintaining human motivation and retention is crucial.
4.4. Technology design
The interface design of AI technology influences user experiences, engagement, and acceptance of AI-driven communication. Interfaces should be user-friendly, promoting seamless information exchange and aligning with human cognitive processes. Since AI systems need time to process human input and generate output, there may be a significant time lag between the interactions in some instances. The delay of feedback between human input and AI responses can adversely affect (MacKenzie and Ware, 1993) the coordination and interaction efficiency between humans and AI, especially in collaborative tasks. On the other hand, human behavior possesses specific temporal patterns, mainly determined by circadian rhythm (Yasseri et al., 2012a). It is essential to design such hybrid systems so that the temporal characteristics of both entities match, support mutual understanding, and balance the cognitive load and workload distribution. Moreover, these systems should be designed to augment human contributions (Jarrahi, 2018).
4.5. Ethical considerations
When integrating AI into human interactions, several ethical considerations emerge. Addressing these concerns involves balancing benefits and minimizing potential harms. Key issues include the ethical implications of mandatory AI use, especially when humans are reluctant or do not require AI assistance. When it comes to AI transparency, a study reported that humans who were paired with bots to play a repeated prisoner’s dilemma game outperformed human-human teams initially but later started to defect once they became aware that they were playing with a bot, showing a trade-off between efficiency and transparency (Crandall et al., 2018). Another critical aspect is addressing biases inherent in AI outputs, which are influenced by the data inputs. It is essential to consider how AI integration might exacerbate or introduce new biases in decision-making (Silberg and Manyika, 2019). Furthermore, liability questions arise when AI errors cause serious harm, an issue compounded by the lack of comprehensive regulations and laws. For instance, a robot failed in object recognition and crushed a human worker to death in a factory in South Korea (BBC News, 2023). This highlights the urgent need for comprehensive legal frameworks to address such liabilities and ensure safety in environments where AI interacts with humans. As AI becomes more prevalent in the workplace, constructing ethical human-AI teams becomes a complex, evolving challenge (Flathmann et al., 2021).
4.6. Conclusion
The advancement of AI technologies has led to enhanced and evolved approaches to tackling existing challenges. Approaches relying solely on human or artificial intelligence encounter significant limitations. Therefore, a synergistic combination of both, complementing each other, is considered the ideal strategy for effective problem-solving.
In current real-world AI-CI applications, AI is predominantly utilized as a technical tool to facilitate data analysis through techniques such as natural language processing and computer vision. Implementing AI not only scales up tasks but, in combination with human intelligence, leads to superior performance. Although AI is critical in some decision-making processes, humans make the ultimate decisions and contributions. It is essential to recognize that AI’s function is to support and enhance human collaborative processes rather than to replace human intelligence. Nevertheless, it is also worth considering collective intelligence emerging from a group of AI entities in future research. Past work has demonstrated that the collective behavior of autonomous machines can lead to the emergence of complex social behavior (Tsvetkova et al., 2017a).
AI-CI approaches present numerous challenges. Success in this field relies on multiple factors beyond the initial idea and readiness of technological solutions. Critical aspects include attracting the audience, designing user-friendly interfaces, and effectively integrating AI with human collective intelligence. Understanding how to combine human intelligence and AI to address social challenges remains a significant and worthy area of study.
To better understand the collectives of humans and AI to help overcome societal challenges, several researchers propose to build the foundations of Social AI (Pedreschi et al., 2023) integrating insights from Complex Systems, Network Science, and AI. We agree that researchers across all relevant disciplines should collaborate and keep up with the latest developments to foster a promising future of AI-enhanced collective intelligence.
Acknowledgements.
The research conducted in this publication was funded by the Irish Research Council under grant number IRCLA/2022/3217, ANNETTE (Artificial Intelligence Enhanced Collective Intelligence). We thank Gianni Giacomelli for useful suggestions and the Supermind Design Augmented Collective Intelligence Database. We thank Siobhan Grayson for the valuable comments on the manuscript.